Financial data for Joel de Paris, Inc., for last year follow: Joel de Paris, Inc. Balance Sheet Beginning Balance Ending Balance Assets Cash $ 137,000 $ 128,000 Accounts receivable 334,000 480,000 Inventory 578,000 480,000 Plant and equipment, net 881,000 882,000 Investment in Buisson, S.A. 398,000 430,000 Land (undeveloped) 247,000 252,000 Total assets $ 2,575,000 $ 2,652,000 Liabilities and Stockholders' Equity Accounts payable $ 384,000 $ 340,000 Long-term debt 984,000 984,000 Stockholders' equity 1,207,000 1,328,000 Total liabilities and stockholders' equity $ 2,575,000 $ 2,652,000 Joel de Paris, Inc. Income Statement Sales $ 5,265,000 Operating expenses 4,422,600 Net operating income 842,400 Interest and taxes: Interest expense $ 110,000 Tax expense 204,000 314,000 Net income $ 528,400 The company paid dividends of $407,400 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%. Required: 1. Compute the company's average operating assets for last year. 2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.) 3. What was the company’s residual income last year?1. Average operating assets Margin Turnover ROI % 3. Residual income
Answer #1

Average operating assets

= [(2,575,000-247,000-398,000)+(2,652,000-252,000-430,000)]/2

= 1,950,000

.

Margin = Net operating income/Sales

= 842,400/5,265,000

= 16%

.

Turnover = Sales/Average operating assets

= 5,265,000/1,950,000

= 2.7

.

ROI = margin * turnover = 16*2.7

= 43.2%

.

Residual income = Net operating income - Required return

= 842,400 - (1,950,000*15%)

= 549,900